OpenStudy (anonymous):

sqrt12a^2 b Simplify by factoring . Assume all expressions under radicals represent non negative numbers. (Type an exact answer, using radicals as needed.)

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{12a^2b}\] \[\Large \sqrt{4*3a^2b}\] \[\Large \sqrt{4}*\sqrt{3}*\sqrt{a^2}*\sqrt{b}\] \[\Large 2*\sqrt{3}*a*\sqrt{b}\] \[\Large 2a\sqrt{3b}\] So \[\Large \sqrt{12a^2b} = 2a\sqrt{3b}\]